### v2

# CDQuant: Accurate Post-training Weight Quantization of Large Pre-trained Models using Greedy Coordinate Descent

Abstract
Large language models (LLMs) Work done when the authors were at Google Research. have recently demonstrated remarkable performance across diverse language tasks. But their deployment is often constrained by their substantial computational and storage requirements. Quantization has emerged as a key technique for addressing this challenge, enabling the compression of large models with minimal impact on performance. The recent GPTQ algorithm, a post-training quantization (PTQ) method, has proven highly effective for compressing LLMs, sparking a wave of research that leverages GPTQ as a core component. Recognizing the pivotal role of GPTQ in the PTQ landscape, we introduce CDQuant, a simple and scalable alternative to GPTQ with improved performance. CDQuant uses coordinate descent to minimize the layer-wise reconstruction loss to achieve high-quality quantized weights. Our algorithm is easy to implement and scales efficiently to models with hundreds of billions of parameters. Through extensive evaluation on the PaLM2 model family, we demonstrate that CDQuant consistently outperforms GPTQ across diverse model sizes and quantization levels. In particular, for INT2 quantization of PaLM2-Otter, CDQuant achieves a $10\%$ reduction in perplexity compared to GPTQ.
1 Introduction
Large language models (LLMs) have shown remarkable ability to handle various language tasks (Touvron et al., 2023; OpenAI, 2023; Google, 2023), but their widespread adoption is hampered by their substantial computational and memory demands. To tackle this challenge, researchers have explored techniques like quantization, pruning, and distillation, with quantization being a particularly promising avenue for reducing model size and inference time without significantly sacrificing performance (Miao et al., 2023). Quantization techniques broadly fall into two categories: post-training quantization (PTQ) and quantization-aware training (QAT). QAT, while potentially yielding better results, poses a significant challenge for LLMs due to the immense resources required to train these large models. As a result, a growing body of research has focused on PTQ, which is generally less computationally intensive and can be applied to large, pre-trained models.
A seminal work in PTQ for LLMs is GPTQ (Frantar et al., 2022), which introduced a one-shot weight quantization approach that minimizes the following layer-wise output reconstruction loss: $\|X(W-\hat{W})\|_{F}^{2}$, where $X=[\mathbf{x}_{1},\mathbf{x}_{2}\dots\mathbf{x}_{n}]$ represents the layer inputs and $W,\widehat{W}\in\mathbb{R}^{d_{\text{in}}\times d_{\text{out}}}$ are the original and quantized weight matrices, respectively. This method has sparked a wave of research in PTQ. Subsequent works have built upon GPTQ, addressing its limitations and improving it’s performance. For instance, SpQR (Dettmers et al., 2023) and OWQ (Lee et al., 2023) improve GPTQ by explicitly addressing outlier weights, maintaining them in full precision, while quantizing the remaining weights using GPTQ. This hybrid approach improves quantization accuracy, particularly at lower bit-widths. Another line of research focuses on transforming the weight space before applying GPTQ. For instance, QuIP (Chee et al., 2024), FrameQuant (Adepu et al., 2024) transform the weights into a more quantization-friendly space and apply GPTQ in the transformed space. AWQ (Lin et al., 2023), SmoothQuant (Xiao et al., 2023) reduce the effect of activation outliers by performing feature scaling before quantizing the weights using standard techniques.
Given the central role GPTQ plays in the landscape of PTQ methods for LLMs, improving its quality directly translates to better quantization across numerous techniques that build upon it. Consequently, in this work, we revisit the core optimization problem addressed by GPTQ and introduce a novel coordinate descent based algorithm, CDQuant, to minimize the objective. CDQuant is an iterative optimization technique that greedily selects coordinates to descend along in each iteration. This approach contrasts with GPTQ, which cycles through the coordinates only once, in a predetermined order, often leading to suboptimal solutions for the layer-wise objective (see Section 2 for more details). CDQuant is simple to implement and scales effectively to models with billions of parameters. We further extend CDQuant  to group/sub-channel quantization, a technique gaining popularity of late (Dettmers et al., 2021; Dettmers and Zettlemoyer, 2023). Extensive experiments on PaLM 2 (Anil et al., 2023) demonstrate that CDQuant consistently outperforms GPTQ across various model sizes and quantization precision levels. For example, for INT2 quantization of PaLM2-Otter, CDQuant achieves a $10\%$ reduction in perplexity compared to GPTQ.
2 Related Work
The literature on quantization is huge. In this section, we only review the works that are related to quantization of large pre-trained models, and those that are related to our work.
Weight-only Quantization.
Optimal Brain Quantization (OBQ) (Frantar and Alistarh, 2022), a post-training quantization (PTQ) technique, draws inspiration from the Optimal Brain Surgeon (OBS) framework proposed by LeCun et al. (1989). Frantar and Alistarh (2022) introduced exact yet efficient algorithms for implementing the OBS greedy strategy. However, OBQ remains computationally expensive and struggles to scale effectively to large language models (LLMs). Frantar et al. (2022) subsequently introduced GPTQ, which employs heuristics to accelerate the OBQ algorithm. Specifically, GPTQ updates coordinates once in a cyclic manner rather than the greedy approach used in OBQ. This modification significantly speeds up the algorithm but comes at the cost of reduced performance. In our work, we address this performance drop by proposing greedy coordinate descent algorithms that are both straightforward and easy to implement.
Recent works have sought to improve upon GPTQ. One popular strategy here is to identify and isolate problematic weights before applying GPTQ. For example, SpQR (Dettmers et al., 2023) isolates outlier weights and keeps them at full precision, quantizing only the remaining values using GPTQ. Similarly, OWQ (Lee et al., 2023) identifies weights that are highly sensitive to small perturbations and excludes them from the GPTQ quantization process. Another promising direction involves transforming the weights into a different basis before quantization. Both QuIP (Chee et al., 2024), and FrameQuant (Adepu et al., 2024) take this route and perform quantization in the transformed space using GPTQ. In all these approaches, we could potentially substitute GPTQ with CDQuant and expect to see further improvements in performance. Other promising methods such as AWQ (Lin et al., 2023), AffineQuant (Ma et al., 2024), suppress the effect of outliers in input activations by adjusting their scale and transferring it to weights. They then perform the standard MinMax quantization on the rescaled weights. One could replace MinMax with GPTQ, CDQuant to improve the performance of these techniques (see Table 3 for a comparison of CDQuant with AWQ).
While working on our paper, we came across QuantEase (Behdin et al., 2023), a parallel research effort sharing similar goals as ours for improving GPTQ. While both methods leverage the concept of coordinate descent, QuantEase adopts a cyclic approach for weight updates, whereas our work employs a greedy strategy. Although our experiments (Appendix F) indicate comparable performance between the two methods, our work extends beyond QuantEase by introducing specialized algorithms for group/sub-channel quantization, not just full-channel quantization. Additionally, we develop novel block coordinate descent algorithms that further improve the performance. That being said, both these works (QuantEase and ours) collectively highlight the potential of coordinate descent algorithms in outperforming GPTQ.
Weight+Activation Quantization.
SmoothQuant (Xiao et al., 2023), OS+(Wei et al., 2023, 2022) have a similar flavour as AWQ, but quantize both weights and activations after scaling them appropriately. OmniQuant (Shao et al., 2023) performs quantization of the entire transformer block in a single shot. This encompasses both activation and weight quantization. Furthermore, it subsumes SmoothQuant, OS+ by using both feature scaling and outlier suppression. QLLM (Liu et al., 2023) tackles the issue of outliers in the activations by splitting the outlier features into multiple sub-channels and then recombining them, effectively reducing their influence. QLLM also incorporates a fine-tuning step at the end, introducing low-rank weights into each layer of the LLM. LLM.int8() (Dettmers et al., 2022) quantizes both acativations and weights to $8$-bits and also identifies outliers and stores them in full precision.
Efficient End-to-End Quantization.
To recover the drop in performance from quantization Chai et al. (2023); Dettmers et al. (2024); Hu et al. (2021) perform low-rank parameter-efficient fine-tuning.
3 CDQuant
Notation.
Throughout the paper, we denote vectors by bold faced letters ($\mathbf{a}$), and matrices by capital letters ($A$). $\|\mathbf{a}\|_{2}=\sqrt{\sum_{i}\mathbf{a}_{i}^{2}}$ is the Euclidean norm and $\|A\|_{F}=\sqrt{\sum_{i,j}A_{ij}^{2}}$ is the Frobenius norm of a matrix. $\mbox{diag}(\mathbf{a})$ represents a diagonal matrix with $\mathbf{a}$ as its diagonal entries. $d_{\text{in}},d_{\text{out}}$ denote the input, output dimensions of a layer. $W\in\mathbb{R}^{d_{\text{in}}\times d_{\text{out}}}$ is the weight matrix of the layer, and $X=[\mathbf{x}_{1},\mathbf{x}_{2}\dots\mathbf{x}_{n}]\in\mathbb{R}^{n\times d_{\text{in}}}$ is the matrix containing $n$ datapoints that are sent as input to the layer. $H=X^{T}X$ is the Hessian matrix for our objective in Equation (1). $c$ denotes the number of bits of precision used in quantization. In quantization, we aim to represent $W$ as $Q\times\mbox{diag}(\mathbf{a})+\mathbf{1}\mathbf{b}^{T}$, where $\mathbf{a},\mathbf{b}\in\mathbb{R}^{d_{\text{out}}}$ represent the scale and bias parameters and $Q\in\{0,1,\dots 2^{c}-1\}^{d_{\text{in}}\times d_{\text{out}}}$ is the quantized matrix.
Many existing PTQ techniques, including GPTQ, aim to solve the following layer-wise optimization objective: $\min_{\mathbf{a},\mathbf{b},Q}\|X(W-Q\times\mbox{diag}(\mathbf{a})-\mathbf{1}\mathbf{b}^{T})\|_{F}^{2}$. Observe that this problem breaks down into $d_{\text{out}}$ independent problems across the output dimension. So, in the sequel, we focus on the following problem of quantizing a $d_{\text{in}}$-dimensional vector
$\min_{a,b,\mathbf{q}}\|X(\mathbf{w}-a\mathbf{q}-b)\|_{2}^{2}.$ (1)
For a fixed $(a,b)$, this problem is called Integer Linear Regression problem. It turns out, finding optimal solutions to this problem is NP-hard (Chrétien and Corset, 2009; Park and Boyd, 2018). So, several works have designed heuristics to solve this problem (Nagel et al., 2020; Li et al., 2021; Frantar and Alistarh, 2022; Frantar et al., 2022; Hubara et al., 2021). Within the context of LLMs, GPTQ is perhaps the most popular among these techniques, as it scales efficiently to models with billions of parameters (Frantar et al., 2022). In this work, we aim to improve upon GPTQ by designing better heuristics, while maintaining its scalability aspect. To this end, we rely on performing coordinate descent on objective (1), which we describe below.
3.1 Greedy Coordinate Descent
Algorithm 1 Greedy Coordinate Descent (CD)
1:Input: $T$ coordinate descent steps, $X$ input data matrix, $\mathbf{w}$ vector to be quantized, $a$ scale, $b$ bias, $\mathbf{q}_{0}$ initial estimate
2:Compute Hessian $H$ as: $H\leftarrow X^{T}X$
3:Compute gradient $\mathbf{g}$ as: $\mathbf{g}\leftarrow 2H(\mathbf{q}_{0}-a^{-1}(\mathbf{w}-b))$
4:for $t\in[1:T]$ do
5:     Find the coordinate that leads to the largest reduction in loss
$i^{*},r^{*}=\operatorname*{arg\,min}_{i\in\{0,1,\dots d_{\text{in}}-1\},r\in\{0,1,\dots 2^{c}-1\}}(r-\mathbf{q}_{t-1,i})^{2}H_{i,i}+(r-\mathbf{q}_{t-1,i})\mathbf{g}_{i}$
6:     Update gradient $\mathbf{g}$ as
$\mathbf{g}\leftarrow\mathbf{g}+2(r^{*}-\mathbf{q}_{t-1,i^{*}})H_{i^{*},\cdot},$
    where $H_{i^{*},\cdot}$ is the $i^{*}$ column of $H$
7:     Update $\mathbf{q}_{t-1}$ as
$\mathbf{q}_{t}\leftarrow\mathbf{q}_{t-1}+(r^{*}-\mathbf{q}_{t-1,i^{*}})\mathbf{e}_{i^{*}},$
   where $\mathbf{e}_{i^{*}}$ is the standard basis vector with $1$ in $i^{*}$ position and 00 everywhere else
8:end for
In this section, we assume we have suitable values for scale ($a$) and bias ($b$) parameters already at hand, and focus on optimizing $\mathbf{q}$. For a more in-depth discussion of how we determine these values, please refer to the final part of the section. As the name suggests, in greedy coordinate descent, at each round, we find the coordinate that leads to the biggest reduction in the objective and descend along that coordinate. Letting ${\mathcal{L}}(\mathbf{q})\coloneqq\|X(\mathbf{w}-a\mathbf{q}-b)\|_{2}^{2}$ be the objective in Equation (1), we try to find a coordinate $i$ and value $r$, such that updating the $i^{th}$ coordinate to $r$ gives the biggest reduction in loss
$\min_{i,r}{\mathcal{L}}(\mathbf{q}+(r-\mathbf{q}_{i})\mathbf{e}_{i})-{\mathcal{L}}(\mathbf{q}),$
where $\mathbf{e}_{i}$ is the standard basis vector with $1$ in $i$ position and 00 everywhere else. Luckily for us, this can be implemented extremely efficiently as we have analytical expressions for the objective. In particular, one can easily show that
${\mathcal{L}}(\mathbf{q}+(r-\mathbf{q}_{i})\mathbf{e}_{i})-{\mathcal{L}}(\mathbf{q})=(r-\mathbf{q}_{i})^{2}H_{i,i}+(r-\mathbf{q}_{i})\mathbf{g}_{i},$
where $H,\mathbf{g}$ are the Hessian and gradient of ${\mathcal{L}}$ evaluated at $\mathbf{q}$. Algorithm 1 describes this procedure. Observe that line 5 is the most expensive step of the algorithm. It requires finding the minimum value among $d_{\text{in}}\times 2^{c}$ possibilities, a task that is well-suited for parallelization on GPUs.
Extension to Block Coordinate Descent.
A natural extension to Algorithm 1 is block coordinate descent (BCD), where multiple coordinates are updated simultaneously in each iteration. While a greedy approach to BCD could, in principle, optimize the objective much better, the computational cost becomes prohibitive. Specifically, updating $k$ coordinates at a time necessitates evaluating $(d_{\text{in}}\times 2^{c})^{k}$ possible combinations of coordinates and their corresponding values. To overcome this, we propose a randomized BCD strategy (Algorithm 2), where we randomly partition the coordinates into $d_{\text{in}}/k$ blocks, and only search over these blocks and their corresponding values. This significantly reduces the search space to a more manageable $d_{\text{in}}/k\times 2^{kc}$ possibilities, making the algorithm practical for larger models. In our experiments, we primarily set $k=2$ and work with $c\in\{2,3,4\}$.
Initializing $a,b,\mathbf{q}$.
To initialize $a,b,\mathbf{q}$ in Algorithms 12, we introduce a technique called Optimal Weight Clipping (OWC), which draws inspiration from the Learnable Weight Clipping (LWC) mechanism used in OmniQuant (Shao et al., 2023). In OWC, we quantize weight $\mathbf{w}$ as follows
$\mathbf{q}=\text{clamp}\left(\left\lfloor\frac{\mathbf{w}-b}{a}\right\rceil,0,2^{c}-1\right),\quad a=\frac{\gamma(\max(\mathbf{w})-\min(\mathbf{w}))}{2^{c}-1},\quad b=\min(\mathbf{w}).$ (2)
Here, $\gamma\in[0,1]$ represents the clipping strength. We determine the optimal $\gamma$ by minimizing the following layer-wise loss:
$\min_{\gamma\in[0,1]}\|X(\mathbf{w}-a\mathbf{q}-b)\|_{2}^{2}.$
Note that while not explicitly stated, both $\mathbf{w},a$ in the above objective are implicitly dependent on $\gamma$. This optimization can be efficiently solved using a simple grid search. In contrast, LWC (Shao et al., 2023) optimizes a different objective function, focusing on end-to-end quantization of entire transformer block using gradient-based techniques. It is worth noting that setting $\gamma=1$ in OWC recovers the widely used MinMax quantization scheme, which is used in many existing quantization methods, including GPTQ (Frantar et al., 2022), SmoothQuant‘(Xiao et al., 2023). In our experiments, we observed that OWC provides a much better initialization compared to MinMax quantization, for both GPTQ and CDQuant.
Algorithm 2 Block Coordinate Descent with Random Blocks (BCD)
1:Input: $T$ coordinate descent steps, $k$ block size, $X$ input data matrix, $\mathbf{w}$ vector to be quantized, $a$ scale, $b$ bias, $\mathbf{q}_{0}$ initial estimate
2:Compute Hessian $H$ as: $H\leftarrow X^{T}X$
3:Compute gradient $\mathbf{g}$ as: $\mathbf{g}\leftarrow 2H(\mathbf{q}_{0}-a^{-1}(\mathbf{w}-b))$
4:for $t\in[1:T]$ do
5:     Randomly partition the set $\{0,1,\dots d_{\text{in}}-1\}$ into $d_{\text{in}}/k$ blocks, each of size $k$
6:     Find the block that leads to the largest reduction in loss
$i^{*},r^{*}=\operatorname*{arg\,min}_{\begin{subarray}{c}i\in\{0,1,\dots d_{\text{in}}/k-1\},\\ r\in\{0,1,\dots 2^{c}-1\}^{k}\end{subarray}}(r-\mathbf{q}_{t-1,i})^{T}H_{i,i}(r-\mathbf{q}_{t-1,i})+(r-\mathbf{q}_{t-1,i})^{T}\mathbf{g}_{i},$
   where $\mathbf{q}_{t-1,i},H_{i,i}$ are the sub-vector, sub-matrix of $\mathbf{q}_{t-1},H$ corresponding to block $i$.
7:     Update gradient $\mathbf{g}$ as
$\mathbf{g}\leftarrow\mathbf{g}+2H_{i^{*},\cdot}(r^{*}-\mathbf{q}_{t-1,i^{*}}),$
8:     Update $\mathbf{q}_{t-1}$ as
$\mathbf{q}_{t-1,i^{*}}\leftarrow r^{*},\quad\mathbf{q}_{t}\leftarrow\mathbf{q}_{t-1},$
9:end for
3.2 Extension to Sub-channel Quantization
In this section, we consider sub-channel (or group) quantization, which is a more fine-grained quantization that divides the weight vector $\mathbf{w}$ into multiple groups and assigns a quantization scale to each group (Dettmers et al., 2021; Dettmers and Zettlemoyer, 2023). Letting $g$ be the group size, we divide weight $\mathbf{w}$ into $d_{\text{in}}/g$ groups $\{\mathbf{w}^{(0)},\mathbf{w}^{(1)},\dots\mathbf{w}^{(d_{\text{in}}/g-1)}\}$ each of size $g$, and quantize $\mathbf{w}^{(i)}$ as $a^{(i)}\mathbf{q}^{(i)}+b^{(i)}$. To learn the optimal parameters for this sub-channel quantization, we solve the following optimization problem:
$\min_{\{a^{(i)},b^{(i)},\mathbf{q}^{(i)}\}_{i=0}^{d_{\text{in}}/g-1}}\big{|}\big{|}\sum_{i=0}X^{(i)}(\mathbf{w}^{(i)}-a^{(i)}\mathbf{q}^{(i)}-b^{(i)})\big{|}\big{|}_{2}^{2}.$ (3)
Here, $X^{(i)}$ represents the columns of $X$ corresponding to the indices within group $i$. To solve this optimization problem, we employ a coordinate descent approach, similar to Algorithms 1 and 2 described earlier. That is, given initial values for the scaling and bias parameters, we iteratively optimize the quantized representation $\mathbf{q}$ using coordinate descent. Due to space constraints, we present the resulting algorithms in Appendix C (see Algorithms 45).
Initialization.
Next, we tackle the initialization of parameters $\{a^{(i)},b^{(i)},\mathbf{q}^{(i)}\}_{i=0}^{d_{\text{in}}/g-1}$ for our coordinate descent procedure. Our approach draws inspiration from the OWC algorithm described above, adapting its core idea to this problem. In essence, we reframe the initialization problem as one of selecting optimal clipping strengths $\{\gamma^{(i)}\}_{i=0}^{d_{\text{in}}/g-1}$ for groups $\{0,\dots d_{\text{in}}/g-1\}$. This leads us to the following problem
$\min_{\gamma^{(0)},\dots\gamma^{(d_{\text{in}}/g-1)}}\big{|}\big{|}\sum_{i=0}X^{(i)}(\mathbf{w}^{(i)}-a^{(i)}\mathbf{q}^{(i)}-b^{(i)})\big{|}\big{|}_{2}^{2},$ (4)
where
$\mathbf{q}^{(i)}=\text{clamp}\left(\left\lfloor\frac{\mathbf{w}^{(i)}-b^{(i)}}{a^{(i)}}\right\rceil,0,2^{c}-1\right),\ a^{(i)}=\frac{\gamma^{(i)}(\max(\mathbf{w}^{(i)})-\min(\mathbf{w}^{(i)}))}{2^{c}-1},\ b^{(i)}=\min(\mathbf{w}^{(i)}).$
We use greedy coordinate descent to optimize Equation (4). In each iteration, we update the $\gamma^{(i)}$ that leads to biggest drop in loss. This procedure, which we call OWC-CD, is described in Algorithm 3.
Algorithm 3 Coordinate Descent for Optimal Weight Clipping (OWC-CD)
1:Input: $T$ coordinate descent steps, $g$ group size, $X$ input data matrix, $\mathbf{w}$ weight vector, $\Gamma$- grid of possible values for clipping strength
2:for $\beta\in\Gamma$ do
3:     for $i\in[0:d_{\text{in}}/g-1]$ do
4:         Compute quantization residual $\Delta(i,\beta)$ for group $i$ with clipping strength $\beta$ as:
$\Delta(i,\beta)\leftarrow\mathbf{w}^{(i)}-a^{(i)}(\beta)\mathbf{q}^{(i)}(\beta)-b^{(i)}(\beta)$
.    where $a^{(i)}(\beta),b^{(i)}(\beta),\mathbf{q}^{(i)}(\beta)$ are as defined in Equation (2).
5:     end for
6:end for
7:Initialize clipping strengths for each group $\gamma^{(0)},\dots\gamma^{(d_{\text{in}}/g-1)}$
8:Compute Hessian $H$ as: $H\leftarrow X^{T}X$
9:$\mathbf{v}_{i}\leftarrow-2H_{i}(a\mathbf{q}+b-w)$    where $H_{i}$ is the sub-matrix of $H$ corresponding to the columns of group $i$.
10:for $t\in[1:T]$ do
11:     Find the group that leads to the largest reduction in loss
$i^{*},\beta^{*}=\operatorname*{arg\,min}_{\begin{subarray}{c}i\in\{0,1,\dots d_{\text{in}}/g-1\},\\ \beta\in\Gamma\end{subarray}}$ $(\Delta(i,\gamma^{(i)})-\Delta(i,\beta))^{T}H_{i,i}(\Delta(i,\gamma^{(i)})-\Delta(i,\beta))$
$+\mathbf{v}_{i}^{T}(\Delta(i,\gamma^{(i)})-\Delta(i,\beta))$
   where $H_{i,i}$ is the block diagonal element of $H$ corresponding to group $i$.
12:     Update $\mathbf{v}\leftarrow\mathbf{v}+2(\Delta(i^{*},\gamma^{(i^{*})})-\Delta(i^{*},\beta^{*}))^{T}H_{i^{*}}$
13:     Update $\gamma^{(i^{*})}\leftarrow\beta^{*}$
14:end for
4 Experiments
Quantization.
In our experiments, we focus on weight-only quantization. We present results for two scenarios: (1) quantizing the FFN layers and (2) quantizing both the FFN and attention weights. In the former setting, all attention layers are quantized to INT8 using MinMax quantization, and low-bit quantization schemes are only applied to the feed-forward layers. It’s worth pointing out that this setting is very relevant in practice because FFN layers account for majority of the latency, and quantizing attention weights yields negligible latency reduction at the expense of model quality (Samaga et al., 2024). We perform both per-channel and sub-channel quantization (with a group size of $128$). We mainly focus on INT2, INT3 and INT4 quantization.
Models.
We use PaLM2 models (Anil et al., 2023) in all our experiments. In particular, we use PaLM2-Gecko, PaLM2-Otter, PaLM2-Bison models. These models vary in size, with PaLM2-Gecko being the smallest, followed by PaLM2-Otter, and finally PaLM2-Bison.
Baselines.
Since our primary goal is to demonstrate that CDQuant gets improved performance over GPTQ, we have chosen it as the primary baseline in most of our experiments. To further demonstrate the our technique can be used as a plug-and-play replacement for GPTQ, we also include comparisons with AWQ (Lin et al., 2023). For all our experiments, we initialize GPTQ with OWC, and run GPTQ for $T=d_{\text{in}}$ steps. To ensure stability and generalization, GPTQ regularizes its Hessian matrix by adding a scaled identity matrix ($\lambda I$). Tuning this $\lambda$ for every (model, layer) pair is infeasible. So, we determine a single optimal value using the PaLM2-Gecko model, and apply it universally.
CDQuant.
We evaluate both the coordinate descent variants described in Section 3.1: CD (Algorithm 1), BCD (Algorithm 2). For per-channel quantization, we initialize CD with Optimal Weight Clipping (OWC), and for sub-channel quantization, we additionally include initialization with OWC-CD. BCD is always initialized with CD in our experiments. Unless otherwise stated, both CD and BCD are run for $T=d_{\text{in}}$ iterations, OWC-CD is run for $d_{\text{in}}/g$ iterations, where $g$ is the group size. Similar to GPTQ, we regularize the Hessian matrix used in CD, BCD by adding $\lambda I$. We determine a reasonable value for $\lambda$ using the PaLM2-Gecko model, and use it in all our experiments.
Evaluation.
Following recent works (Frantar et al., 2022; Ma et al., 2024), we evaluate all algorithms using two key metrics: perplexity and downstream task performance. To ensure a robust evaluation, we calculate perplexity using a $100$ million token subset derived from the PaLM2 training mixture. We chose to do this since the evaluation sets for PTB (Marcus et al., 1994) and WikiText2 (Merity et al., 2017) datasets, that are commonly used in the field, are orders of magnitude smaller (83K and 246K tokens in their respective test sets, and 512K tokens subsampled from the C4 validation set  (Frantar et al., 2022)), and thus can have high variance, and may not accurately reflect the model’s true performance. We use TriviaQA (Joshi et al., 2017), SQuAD (Rajpurkar et al., 2018), NaturalQuestions (Kwiatkowski et al., 2019) and WebQuestions (Berant et al., 2013) to evaluate generation capabilities of the quantized models, and to evaluate their reasoning capabilities, we test on ARC-c, ARC-e (Clark et al., 2018), HellaSwag (Zellers et al., 2019), BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2020) and WinoGrande (Sakaguchi et al., 2020). For downstream evaluations, we consider both zero-shot, and one-shot settings. Finally, to determine which optimization technique is most effective at solving the layer-wise objective in Equation (1), we evaluate the final solution’s objective value using the same $100$ million tokens that we used to compute perplexity.
Training.
All techniques are calibrated using $1280$ data points, where each data point has $2048$ tokens. For OWC, we use a grid size of $50$ to find the most optimal $\gamma$. We used 8 Nvidia H100100100100 GPUs for quantizing the models.
Table 1: Table presents pretraining perplexity numbers for GPTQ, CD, BCD for INT3, INT4 quantization of FFN weights. $wx,ay,gz$ in the config column corresponds to $x$-bit weights, $y$-bit activations and group/sub-channel size of $z$.
Config Method PaLM2-Gecko PaLM2-Otter PaLM2-Bison
w16a16 - 7.948 5.984 5.298
w3a16 OWC $12.570$ $17.928$ $6.169$
GPTQ $11.347$ $7.176$ $5.774$
CD $10.920$ $7.002$ $5.739$
BCD(k=2) $\mathbf{10.898}$ $\mathbf{6.979}$ $\mathbf{5.733}$
w3a16g128 OWC $11.597$ $8.342$ $5.847$
GPTQ $10.414$ $6.635$ $5.677$
CD $10.273$ $6.655$ $5.656$
BCD(k=2) $10.259$ $6.545$ $5.654$
OWC-CD $10.706$ $6.635$ $5.686$
OWC-CD + CD $10.143$ $6.528$ $5.650$
OWC-CD + BCD(k=2) $\mathbf{10.138}$ $\mathbf{6.527}$ $\mathbf{5.647}$
w4a16 OWC $8.946$ $6.693$ $5.475$
GPTQ $8.764$ $6.249$ $5.417$
CD $8.694$ $6.195$ $5.407$
BCD(k=2) $\mathbf{8.691}$ $\mathbf{6.192}$ $\mathbf{5.405}$
w4a16g128 OWC $8.613$ $6.264$ $5.401$
GPTQ $8.498$ $6.112$ $5.377$
CD $8.456$ $6.097$ $5.373$
BCD(k=2) $8.454$ $6.097$ $5.372$
OWC-CD $8.519$ $6.106$ $5.377$
OWC-CD + CD $8.436$ $6.092$ $5.371$
OWC-CD + BCD(k=2) $\mathbf{8.434}$ $\mathbf{6.091}$ $\mathbf{5.371}$
Table 2: Table presents downstream evaluation (one-shot) numbers for GPTQ, CD, BCD for INT3, INT4 quantization of FFN weights. Gen, Rank columns correspond to generation and ranking tasks.
Config Method PaLM2-Gecko PaLM2-Otter PaLM2-Bison
Gen. Rank Avg. Gen. Rank Avg. Gen. Rank Avg.
w16a16 - $25.85$ $61.44$ $47.2$ $44.31$ $75.43$ $62.98$ $51.08$ $78.28$ $68.00$
w3a16 OWC $18.95$ $55.74$ $41.02$ $21.02$ $65.78$ $47.87$ $47.52$ $77.76$ $65.66$
GPTQ $19.42$ $57.71$ $42.39$ $37.26$ $73.95$ $59.27$ $48.55$ $78.58$ $66.57$
CD $20.69$ $58.05$ $43.11$ $38.12$ $73.37$ $59.27$ $48.47$ $78.56$ $66.53$
BCD(k=2) $20.52$ $58.44$ $43.27$ $38.53$ $73.46$ $59.49$ $48.08$ $78.66$ $66.43$
w3a16g128 OWC $22.83$ $57.02$ $43.35$ $33.18$ $70.81$ $55.76$ $48.08$ $78.34$ $66.24$
GPTQ $22.24$ $58.66$ $44.09$ $40.30$ $74.12$ $60.59$ $49.15$ $78.58$ $66.81$
CD $22.79$ $58.81$ $44.40$ $41.33$ $74.27$ $61.10$ $49.17$ $78.49$ $66.76$
BCD(k=2) $22.64$ $58.85$ $44.37$ $41.19$ $73.94$ $60.84$ $49.02$ $78.51$ $66.71$
OWC-CD $22.21$ $58.69$ $44.10$ $40.84$ $74.33$ $60.93$ $48.54$ $78.29$ $66.39$
OWC-CD + CD $22.44$ $58.84$ $44.28$ $40.87$ $74.41$ $61.00$ $49.22$ $78.86$ $67.00$
OWC-CD + BCD(k=2) $22.15$ $58.77$ $44.13$ $41.19$ $74.65$ $61.27$ $49.08$ $78.3$ $66.61$
w4a16 OWC $23.27$ $60.22$ $45.44$ $41.70$ $74.21$ $61.21$ $50.05$ $78.72$ $67.25$
GPTQ $24.51$ $60.72$ $46.24$ $43.00$ $74.86$ $62.12$ $50.28$ $79.15$ $67.6$
CD $24.60$ $60.62$ $46.22$ $43.21$ $75.17$ $62.39$ $50.09$ $79.21$ $67.56$
BCD(k=2) $24.52$ $60.62$ $46.18$ $43.16$ $75.50$ $62.56$ $50.25$ $79.14$ $67.59$
w4a16g128 OWC $24.63$ $61.27$ $46.61$ $42.31$ $74.76$ $61.78$ $50.49$ $79.38$ $67.82$
GPTQ $25.40$ $61.13$ $46.84$ $43.54$ $75.02$ $62.43$ $50.63$ $79.21$ $67.78$
CD $25.18$ $61.42$ $46.92$ $43.78$ $74.93$ $62.47$ $50.73$ $79.30$ $67.87$
BCD(k=2) $24.95$ $61.49$ $46.87$ $43.77$ $75.04$ $62.53$ $50.84$ $79.56$ $68.07$
OWC-CD $24.59$ $61.40$ $46.67$ $43.67$ $75.31$ $62.65$ $50.65$ $79.04$ $67.69$
OWC-CD + CD $25.12$ $61.44$ $46.91$ $44.00$ $75.27$ $62.76$ $50.62$ $79.21$ $67.78$
OWC-CD + BCD(k=2) $25.00$ $61.50$ $46.90$ $43.75$ $75.20$ $62.62$ $50.64$ $79.17$ $67.76$
Results.
Table 1 presents the perplexity numbers for different quantization techniques applied to FFN layers. It can be seen that both our CD and BCD methods have a clear advantage over GPTQ, leading to lower perplexity scores for all models and quantization levels. The difference is more pronounced for lower bit quantization. This is further highlighted in Table 5, which focuses specifically on INT2 quantization. Here, we see almost $10\%$ improvement in perplexity over GPTQ. We also present (1-shot) downstream evals in Tables 25 (for 0-shot results, and a detailed breakdown of results for each dataset, please refer to Appendix D). In Table 5 we see $2$-$2.5\%$ relative improvement in downstream evals over GPTQ. Furthermore, Table 2 demonstrates that our methods consistently match or exceed GPTQ’s performance across all the settings, with the most substantial improvements observed in low-bit quantization, and smaller models. Finally, we also observe that coordinate descent techniques are better at optimizing the layer-wise objective in Equation (1), than GPTQ. For instance, the average objective value (relative to all 00’s solution) for the GPTQ solution for the $1^{\text{st}}$ feed-forward layer is $0.164$, whereas for CD it is $0.158$, and for BCD(k=$2$) it is $0.157$.
Next, observe that Table 5 also investigates the effect of multiple epochs of CD and BCD, where each epoch corresponds to $d_{\text{in}}$ iterations. For BCD with $k=2$, additional epochs enhance performance. However, for BCD with $k=4$, multiple epochs appear to lead to overfitting, indicating a need for stronger regularization.
FFN+Attention Quantization. In addition to FFN quantization, we also perform FFN + attention weight quantization. However, the inputs to the attention layer are often aligned in a handful of directions. Consequently, performing quantization using such a data leads to a huge drop in performance, as the algorithms would primarily focus on a few directions and ignore the rest. To mitigate this, we clip the largest eigenvalues of the Hessian to ensure a more balanced Hessian. This technique, reminiscent of the weight clipping in OmniQuant (Shao et al., 2023), improves the performance of both GPTQ and CDQuant [^1] One could also rely on existing techniques such as AWQ and SmoothQuant to reduce the effect of outliers. But in our experiments we noticed that both AWQ and SmoothQuant performed poorly compared to the simple eigenvalue clipping technique.. For instance, for PaLM2-Gecko, the perplexity for w3a16-GPTQ improves from $34.872$ to $24.054$ with clipping, and for w4a16-GPTQ, it improves from $10.966$ to $10.005$. Table 4 presents the results from this experiment, where we run all the algorithms on the clipped Hessian. We once again notice that in almost all the settings, our coordinate descent algorithms outperform GPTQ.
Table 3: Perplexity numbers for GPTQ, CD, BCD for quantization of FFN layers of PaLM2-Otter, with and without (w/o) AWQ.
Method PaLM2-Otter
w3a16 w3a16g128
with AWQ w/o AWQ with AWQ w/o AWQ
OWC $22.661$ $17.928$ $6.813$ $8.342$
GPTQ $7.381$ $7.176$ $6.519$ $6.635$
CD $7.244$ $7.002$ $6.475$ $6.655$
BCD(k=2) $\mathbf{7.211}$ $\mathbf{6.979}$ $\mathbf{6.474}$ $\mathbf{6.545}$
AWQ. A major strength of our algorithm is its versatility. It seamlessly replaces GPTQ in any quantization technique that relies on it. To illustrate this, we focus on the AWQ algorithm. We demonstrate that our algorithm, when layered on top of AWQ, surpasses the performance of GPTQ layered on top of AWQ. Table 3 presents the result from this experiment. It can be seen that BCD always outperforms GPTQ, for both per-channel and group quantization  [^2] The performance of AWQ on PaLM2-Bison is much worse than not performing AWQ. Hence, in Table 3, we only present results for PaLM2-Otter..
Runtime.
The runtimes of various techniques for quantizing PaLM2-Otter and PaLM2-Bison are presented in Table 6. To be precise, the table presents time taken to quantize all the FFN, attention layers in the model, for per-channel quantization. For group quantization runtimes, please refer to Appendix E. It can be seen that CD and GPTQ have comparable runtimes. While BCD takes longer than CD, the majority of its time is spent quantizing the $2^{nd}$ FFN layer, where the quantization dimension is equivalent to the model’s hidden dimension. So, in practice, to speed up the entire process, one could rely on CD for quantizing FFN2 and use BCD for quantizing the rest of the weight matrices.
Table 4: Table presents pretraining perplexity numbers for GPTQ, CD, BCD for INT3, INT4 quantization of both FFN and attention weights.
Config Method PaLM2-Gecko PaLM2-Otter PaLM2-Bison
w16a16 - 7.948 5.984 5.298
w3a16 OWC $41.751$ $54.499$ $6.446$
GPTQ $24.054$ $\mathbf{11.384}$ $5.801$
CD $19.692$ $11.542$ $5.770$
BCD(k=2) $\mathbf{19.312}$ $11.392$ $\mathbf{5.762}$
w3a16g128 OWC $20.301$ $9.785$ $5.973$
GPTQ $15.042$ $7.501$ $5.698$
CD $14.757$ $7.454$ $5.679$
BCD(k=2) $14.691$ $\mathbf{7.438}$ $5.673$
OWC-CD $16.268$ $7.617$ $5.746$
OWC-CD + CD $14.285$ $7.466$ $5.675$
OWC-CD + BCD(k=2) $\mathbf{14.242}$ $7.458$ $\mathbf{5.670}$
w4a16 OWC $10.683$ $7.194$ $5.508$
GPTQ $10.005$ $6.645$ $5.425$
CD $9.991$ $6.627$ $5.415$
BCD(k=2) $\mathbf{9.965}$ $\mathbf{6.621}$ $\mathbf{5.413}$
w4a16g128 OWC $9.402$ $6.415$ $5.420$
GPTQ $9.160$ $6.241$ $5.381$
CD $9.124$ $6.226$ $5.377$
BCD(k=2) $9.120$ $\mathbf{6.225}$ $5.376$
OWC-CD $9.241$ $6.261$ $5.387$
OWC-CD + CD $9.094$ $6.244$ $5.375$
OWC-CD + BCD(k=2) $\mathbf{9.093}$ $6.242$ $\mathbf{5.374}$
Table 5: Table presents the perplexity, and downstream evaluations (1-shot) for GPTQ, CD, BCD for INT2 quantization of FFN weights. Gen, Rank columns correspond to generation and ranking tasks.
Method Epochs PaLM2-Otter PaLM2-Bison
w2a16g128 Perplexity Generation Rank Avg. Perplexity Generation Rank Avg.
GPTQ - $10.816$ $26.98$ $67.14$ $51.07$ $7.230$ $39.32$ $74.19$ $60.24$
CD $1$ $9.917$ $29.87$ $66.76$ $52.00$ $7.123$ $38.73$ $73.56$ $59.63$
BCD(k=2) $1$ $9.822$ $30.03$ $66.82$ $52.11$ $7.094$ $38.85$ $73.99$ $59.93$
$2$ $9.760$ $30.09$ $\mathbf{67.45}$ $\mathbf{52.51}$ $\mathbf{7.088}$ $\mathbf{39.35}$ $74.10$ $60.20$
$3$ $9.719$ $\mathbf{30.19}$ $67.09$ $52.33$ $7.094$ $39.30$ $\mathbf{74.24}$ $\mathbf{60.26}$
BCD(k=4) $1$ $\mathbf{9.708}$ $30.11$ $67.09$ $52.30$ $7.099$ $38.98$ $74.20$ $60.11$
$2$ $9.752$ $30.00$ $66.70$ $52.02$ $7.099$ $39.09$ $74.08$ $60.08$
$3$ $9.749$ $30.07$ $66.93$ $52.19$ $7.099$ $39.05$ $74.10$ $60.08$
Table 6: Table presents the per-channel quantization runtime (in minutes) of CD, BCD, GPTQ (each run for $d_{\text{in}}$ iterations) on PaLM2-Otter and PaLM2-Bison.
Config Method PaLM2-Otter PaLM2-Bison
FFN runtime
(in minutes)
Attn. runtime
(in minutes)
FFN runtime
(in minutes)
Attn. runtime
(in minutes)
w3a16 GPTQ $0.90$m $0.72$m $6.42$m $0.63$m
CD $2.94$m $0.49$m $16.13$m $0.43$m
BCD(k=2) $14.03$m $2.08$m $82.13$m $1.81$m
w4a16 GPTQ $0.90$m $0.64$m $6.60$m $0.56$m
CD $3.59$m $0.59$m $21.19$m $0.52$m
BCD(k=2) $34.85$m $4.72$m $235.14$m $4.13$m
5 Conclusion and Future Work
In this work, we developed a coordinate descent framework (CDQuant) for quantization of LLMs. CDQuant is a simple and effective alternative to GPTQ, that consistently outperformed it on PaLM2 models. The simplicity of our algorithm makes it a seamless substitute for GPTQ in various algorithmic contexts where GPTQ currently functions as a sub-routine. Our future work aims to further improve the performance of CDQuant. In particular, we aim to speed up our BCD algorithm, and make it as fast as CD. Furthermore, we will focus on developing layer-wise loss functions that are more closely aligned with end-to-end loss, thereby reducing the performance gap between full-precision and quantized models. Finally, we will also explore integrating CDQuant with QuIP and FrameQuant to evaluate potential performance gains for extreme low-bit quantization.
Acknowledgements
We are grateful to Kyuyeun Kim, Wonpyo Park, and Jaehong Kim for assisting in setting up the calibration pipelines, and Prateek Jain for helpful discussions, support, and feedback.
Appendix A Broader Impact
We introduce a Coordinate Descent based approach, CDQuant, for compressing Large Language Models. Our method uses only a small amount of data for calibration. We do not foresee any ethical implications arising from the technical aspects of our approach. However, compressing LLMs may give rise to bias effects, a study of which seems essential given the extensive use of LLMs. Our work may be of assistance to such studies. Also, since quantization allows for easier deployment of LLMs, it could have potential societal implications which seem difficult to predict.
Appendix B Limitations
  • While both CD and BCD outperformed GPTQ in our experiments, BCD achieves slightly better performance than CD. However, BCD is not as fast as CD and can be expensive for large models. In future, we aim to develop techniques to speed up BCD.
  • Our algorithms still don’t bridge the gap between QAT and PTQ, especially on smaller models. To bridge this gap, we believe one should move away from the $\ell_{2}^{2}$ surrogate loss that is being considered by most of the existing work. Instead, we should design surrogate losses that are more closely aligned with end-to-end loss.
  • Due to computational reasons, we couldn’t run experiments where we replace GPTQ component with CDQuant in algorithms such as QuIP, FrameQuant.
Appendix C CDQuant for sub-channel quantization
To simplify the explanation in this section, we introduce a slightly modified notation. Given a weight vector $\mathbf{w}$, we represent it’s sub-channel quantization as $\mathbf{a}\odot\mathbf{q}+\mathbf{b}$, where $\odot$ is the elementwise multiplication, and $\mathbf{a},\mathbf{b}\in\mathbb{R}^{d_{\text{in}}}$ are the scale, and bias parameters that satisfy the following constraints: $\mathbf{a}_{k}=\mathbf{a}_{l},\mathbf{b}_{k}=\mathbf{b}_{l}$, for any two indices $k,l$ that fall in the same group. With this notation, for any given $\mathbf{a},\mathbf{b}$, the optimization problem in Equation (3) can be rewritten as
$\min_{\mathbf{q}}\|X(\mathbf{w}-\mathbf{a}\odot\mathbf{q}-\mathbf{b})\|_{2}^{2}.$ (5)
Letting $D_{\mathbf{a}}=\mbox{diag}(\mathbf{a}),\tilde{X}=XD_{\mathbf{a}},\tilde{\mathbf{w}}=D_{\mathbf{a}}^{-1}\mathbf{w},\tilde{\mathbf{b}}=D_{\mathbf{a}}^{-1}\mathbf{b}$, the above problem can be further rewritten as
$\min_{\mathbf{q}}\|\tilde{X}(\tilde{\mathbf{w}}-\mathbf{q}-\tilde{\mathbf{b}})\|_{2}^{2}.$ (6)
Observe that this problem is the same as the per-channel quantization problem described in Equation (1), but with modified parameters $\tilde{X},\tilde{\mathbf{w}},\tilde{\mathbf{b}}$. So extending CDQuant to sub-channel quantization simply involves running Algorithms 12 with these modified parameters. Algorithms 45 present the resulting algorithms.
Algorithm 4 Greedy Coordinate Descent (CD)
1:Input: $T$ coordinate descent steps, $X$ input data matrix, $\mathbf{w}$ vector to be quantized, $a$ scale, $b$ bias, $\mathbf{q}_{0}$ initial estimate
2:$\tilde{X}\leftarrow X\mbox{diag}(\mathbf{a}),\tilde{\mathbf{w}}=\mbox{diag}(\mathbf{a})^{-1}\mathbf{w},\tilde{\mathbf{b}}=\mbox{diag}(\mathbf{a})\mathbf{b}$
3:Compute Hessian $H$ as: $H\leftarrow\tilde{X}^{T}\tilde{X}$
4:Compute gradient $\mathbf{g}$ as: $\mathbf{g}\leftarrow 2H(\mathbf{q}_{0}-(\tilde{\mathbf{w}}-\tilde{\mathbf{b}}))$
5:for $t\in[1:T]$ do
6:     Find the coordinate that leads to the largest reduction in loss
$i^{*},r^{*}=\operatorname*{arg\,min}_{i\in\{0,1,\dots d_{\text{in}}-1\},r\in\{0,1,\dots 2^{c}-1\}}(r-\mathbf{q}_{t-1,i})^{2}H_{i,i}+(r-\mathbf{q}_{t-1,i})\mathbf{g}_{i}$
7:     Update gradient $\mathbf{g}$ as
$\mathbf{g}\leftarrow\mathbf{g}+2(r^{*}-\mathbf{q}_{t-1,i^{*}})H_{i^{*},\cdot},$
    where $H_{i^{*},\cdot}$ is the $i^{*}$ column of $H$
8:     Update $\mathbf{q}_{t-1}$ as
$\mathbf{q}_{t}\leftarrow\mathbf{q}_{t-1}+(r^{*}-\mathbf{q}_{t-1,i^{*}})\mathbf{e}_{i^{*}},$
   where $\mathbf{e}_{i^{*}}$ is the standard basis vector with $1$ in $i^{*}$ position and 00 everywhere else
9:end for
Algorithm 5 Block Coordinate Descent with Random Blocks (BCD)
1:Input: $T$ coordinate descent steps, $k$ block size, $X$ input data matrix, $\mathbf{w}$ vector to be quantized, $a$ scale, $b$ bias, $\mathbf{q}_{0}$ initial estimate
2:$\tilde{X}\leftarrow X\mbox{diag}(\mathbf{a}),\tilde{\mathbf{w}}=\mbox{diag}(\mathbf{a})^{-1}\mathbf{w},\tilde{\mathbf{b}}=\mbox{diag}(\mathbf{a})\mathbf{b}$
3:Compute Hessian $H$ as: $H\leftarrow\tilde{X}^{T}\tilde{X}$
4:Compute gradient $\mathbf{g}$ as: $\mathbf{g}\leftarrow 2H(\mathbf{q}_{0}-(\tilde{\mathbf{w}}-\tilde{\mathbf{b}}))$
5:for $t\in[1:T]$ do
6:     Randomly partition the set $\{0,1,\dots d_{\text{in}}-1\}$ into $d_{\text{in}}/k$ blocks, each of size $k$
7:     Find the block that leads to the largest reduction in loss
$i^{*},r^{*}=\operatorname*{arg\,min}_{\begin{subarray}{c}i\in\{0,1,\dots d_{\text{in}}/k-1\},\\ r\in\{0,1,\dots 2^{c}-1\}^{k}\end{subarray}}(r-\mathbf{q}_{t-1,i})^{T}H_{i,i}(r-\mathbf{q}_{t-1,i})+(r-\mathbf{q}_{t-1,i})^{T}\mathbf{g}_{i},$
   where $\mathbf{q}_{t-1,i},H_{i,i}$ are the sub-vector, sub-matrix of $\mathbf{q}_{t-1},H$ corresponding to block $i$.
8:     Update gradient $\mathbf{g}$ as
$\mathbf{g}\leftarrow\mathbf{g}+2H_{i^{*},\cdot}(r^{*}-\mathbf{q}_{t-1,i^{*}}),$
9:     Update $\mathbf{q}_{t-1}$ as
$\mathbf{q}_{t-1,i^{*}}\leftarrow r^{*},\quad\mathbf{q}_{t}\leftarrow\mathbf{q}_{t-1},$
10:end for
Appendix D Detailed Downstream Evaluation Results
Table 7: Table presents downstream evaluation (one-shot) numbers of PaLM2-Bison - for GPTQ, CD, BCD for INT3, INT4 quantization of FFN weights.
Config Method PaLM2-Bison
NatualQ. SQuAD TriviaQA WebQ ARC-c ARC-e BoolQ HellaSwag PIQA WinoGrande
w16a16 - $27.84$ $75.19$ $77.42$ $23.87$ $59.04$ $84.64$ $88.87$ $82.50$ $82.43$ $78.22$
w3a16 OWC $22.16$ $75.84$ $70.31$ $21.75$ $57.34$ $83.50$ $86.94$ $79.60$ $81.12$ $78.06$
GPTQ $24.29$ $74.65$ $72.29$ $22.98$ $58.70$ $85.02$ $87.58$ $80.81$ $81.72$ $77.66$
CD $23.60$ $74.92$ $72.25$ $23.13$ $58.53$ $84.68$ $88.20$ $80.60$ $81.61$ $77.74$
BCD(k=2) $23.24$ $74.78$ $72.06$ $22.24$ $58.62$ $84.55$ $88.04$ $80.69$ $81.83$ $78.22$
w3a16g128 OWC $23.88$ $73.59$ $73.31$ $21.56$ $57.85$ $83.84$ $88.23$ $80.67$ $80.79$ $78.69$
GPTQ $25.32$ $73.96$ $74.10$ $23.23$ $58.02$ $84.39$ $88.50$ $80.67$ $81.83$ $78.06$
CD $25.24$ $74.16$ $74.11$ $23.18$ $57.51$ $83.75$ $88.38$ $81.15$ $82.15$ $77.98$
BCD(k=2) $24.93$ $74.05$ $73.90$ $23.18$ $56.74$ $84.05$ $88.26$ $81.06$ $82.43$ $78.53$
OWC-CD $24.60$ $74.03$ $73.80$ $21.75$ $57.17$ $84.18$ $88.2$ $80.97$ $81.94$ $77.27$
OWC-CD + CD $25.12$ $74.37$ $73.99$ $23.38$ $58.36$ $85.44$ $88.65$ $80.93$ $81.77$ $77.98$
OWC-CD + BCD(k=2) $25.32$ $73.54$ $74.30$ $23.18$ $57.34$ $83.96$ $88.59$ $81.01$ $81.39$ $77.51$
w4a16 OWC $26.57$ $74.66$ $75.84$ $23.13$ $58.28$ $83.42$ $88.75$ $81.39$ $81.66$ $78.85$
GPTQ $27.06$ $74.87$ $76.45$ $22.74$ $58.87$ $84.60$ $88.53$ $81.72$ $82.48$ $78.69$
CD $26.79$ $74.40$ $76.35$ $22.83$ $59.22$ $84.68$ $88.99$ $81.60$ $82.10$ $78.69$
BCD(k=2) $27.06$ $74.54$ $76.28$ $23.13$ $59.81$ $84.76$ $88.56$ $81.70$ $82.26$ $77.74$
w4a16g128 OWC $27.04$ $75.37$ $76.42$ $23.13$ $59.73$ $84.97$ $88.87$ $81.94$ $82.32$ $78.45$
GPTQ $27.12$ $75.63$ $76.42$ $23.38$ $58.45$ $84.76$ $89.05$ $82.01$ $82.15$ $78.85$
CD $27.31$ $75.52$ $76.45$ $23.62$ $59.04$ $85.02$ $88.50$ $82.04$ $82.54$ $78.69$
BCD(k=2) $27.37$ $75.44$ $76.92$ $23.62$ $58.87$ $85.35$ $88.72$ $81.98$ $82.97$ $79.48$
OWC-CD $27.31$ $75.70$ $76.64$ $22.93$ $58.36$ $84.64$ $88.56$ $81.98$ $82.59$ $78.14$
OWC-CD + CD $27.01$ $75.52$ $76.69$ $23.28$ $58.79$ $85.10$ $88.56$ $81.98$ $82.15$ $78.69$
OWC-CD + BCD(k=2) $26.98$ $75.42$ $76.84$ $23.33$ $58.79$ $85.35$ $88.47$ $81.93$ $82.32$ $78.14$
Table 8: Table presents downstream evaluation (one-shot) numbers of PaLM2-Otter - for GPTQ, CD, BCD for INT3, INT4 quantization of FFN weights.
Config Method PaLM2-Otter
NatualQ. SQuAD TriviaQA WebQ ARC-c ARC-e BoolQ HellaSwag PIQA WinoGrande
w16a16 - $19.78$ $70.3$ $69.1$ $18.06$ $54.1$ $80.13$ $85.47$ $79.51$ $79.43$ $73.95$
w3a16 OWC $5.96$ $50.17$ $24.8$ $3.15$ $45.99$ $66.84$ $74.83$ $66.98$ $72.47$ $67.56$
GPTQ $13.77$ $66.12$ $56.15$ $12.99$ $52.56$ $78.58$ $85.17$ $76.75$ $78.35$ $72.3$
CD $14.49$ $65.71$ $58.83$ $13.44$ $52.47$ $79.92$ $83.03$ $75.81$ $77.09$ $71.9$
BCD(k=2) $14.63$ $65.96$ $59.46$ $14.07$ $52.73$ $79.84$ $83.09$ $75.93$ $77.42$ $71.74$
w3a16g128 OWC $10.89$ $67.98$ $45.03$ $8.81$ $48.29$ $74.83$ $79.14$ $74.64$ $75.52$ $72.45$
GPTQ $16.12$ $69.8$ $60.82$ $14.47$ $53.16$ $79.34$ $84.07$ $77.12$ $77.86$ $73.16$
CD $15.46$ $71.62$ $62.07$ $16.19$ $52.3$ $78.79$ $84.34$ $77.02$ $78.35$ $74.82$
BCD(k=2) $16.04$ $71.04$ $62.02$ $15.65$ $51.71$ $78.32$ $84.83$ $77.04$ $77.69$ $74.03$
OWC-CD $15.62$ $71.78$ $61.19$ $14.76$ $52.99$ $79.34$ $84.1$ $76.92$ $77.42$ $75.22$
OWC-CD + CD $16.2$ $70.77$ $61.8$ $14.71$ $53.41$ $79.42$ $84.5$ $77.22$ $78.29$ $73.64$
OWC-CD + BCD(k=2) $16.7$ $71.3$ $61.87$ $14.91$ $53.84$ $79.8$ $84.37$ $77.25$ $78.45$ $74.19$
w4a16 OWC $16.81$ $71.7$ $61.75$ $16.54$ $53.33$ $78.24$ $85.17$ $77.32$ $79.33$ $71.9$
GPTQ $19$ $70.66$ $66.01$ $16.34$ $53.67$ $79.46$ $85.57$ $78.5$ $79.05$ $72.93$
CD $18.7$ $70.97$ $66.11$ $17.08$ $54.27$ $80.89$ $85.87$ $78.55$ $79.05$ $72.38$
BCD(k=2) $19.14$ $70.9$ $65.72$ $16.88$ $54.78$ $80.6$ $85.6$ $78.64$ $79.65$ $73.72$
w4a16g128 OWC $18.03$ $69.97$ $65.28$ $15.94$ $53.41$ $79.42$ $84.89$ $78.1$ $79.16$ $73.56$
GPTQ $19.17$ $70.34$ $67.46$ $17.18$ $53.58$ $79.88$ $85.02$ $79$ $79.11$ $73.56$
CD $19.25$ $71.12$ $67.61$ $17.13$ $53.33$ $80.18$ $84.92$ $79.09$ $79.11$ $72.93$
BCD(k=2) $19.45$ $70.81$ $67.75$ $17.08$ $54.27$ $80.3$ $84.74$ $79.05$ $78.94$ $72.93$
OWC-CD $19.2$ $70.58$ $67.66$ $17.22$ $53.24$ $80.51$ $85.05$ $78.84$ $79.49$ $74.74$
OWC-CD + CD $19.64$ $70.8$ $68.02$ $17.52$ $53.75$ $80.77$ $84.98$ $79.09$ $78.78$ $74.27$
OWC-CD + BCD(k=2) $19.31$ $70.41$ $67.91$ $17.37$ $53.67$ $80.3$ $84.95$ $79.07$ $78.78$ $74.43$
Table 9: Table presents downstream evaluation (one-shot) numbers of PaLM2-Gecko - for GPTQ, CD, BCD for INT3, INT4 quantization of FFN weights.
Config Method PaLM2-Gecko
NatualQ. SQuAD TriviaQA WebQ ARC-c ARC-e BoolQ HellaSwag PIQA WinoGrande
w16a16 - $7.67$ $49.04$ $37.72$ $8.96$ $37.12$ $68.1$ $66.15$ $61.55$ $73.99$ $61.72$
w3a16 OWC $3.71$ $48.17$ $18.84$ $5.07$ $30.03$ $60.1$ $59.69$ $53.45$ $70.24$ $60.93$
GPTQ $4.38$ $46.91$ $20.43$ $5.95$ $31.23$ $62.84$ $63.55$ $56.51$ $71.65$ $60.46$
CD $4.63$ $48.1$ $22.79$ $7.23$ $32.59$ $63.51$ $62.14$ $57.6$ $72.09$ $60.38$
BCD(k=2) $4.52$ $48.84$ $22.09$ $6.64$ $32.51$ $64.14$ $63.24$ $57.53$ $72.2$ $61.01$
w3a16g128 OWC $4.43$ $57.48$ $20.99$ $8.42$ $32.59$ $63.09$ $62.05$ $53.45$ $71.38$ $59.59$
GPTQ $4.85$ $53.14$ $23.25$ $7.73$ $33.11$ $64.73$ $64.43$ $56.19$ $72.42$ $61.09$
CD $4.49$ $54.43$ $24.12$ $8.12$ $33.28$ $65.15$ $64.22$ $56.56$ $71.98$ $61.64$
BCD(k=2) $4.29$ $53.58$ $24.41$ $8.27$ $33.7$ $65.32$ $64.25$ $56.25$ $71.71$ $61.88$
OWC-CD $4.6$ $53.86$ $23.23$ $7.14$ $32.85$ $64.81$ $64.65$ $56.31$ $71.82$ $61.72$
OWC-CD + CD $4.52$ $52.4$ $25.07$ $7.78$ $32.76$ $64.44$ $65.11$ $56.85$ $71.93$ $61.96$
OWC-CD + BCD(k=2) $4.65$ $51.15$ $25.13$ $7.68$ $33.19$ $64.9$ $64.68$ $56.88$ $71.82$ $61.17$
w4a16 OWC $6.09$ $48.4$ $31.06$ $7.53$ $35.24$ $65.07$ $64.59$ $59.59$ $72.96$ $63.85$
GPTQ $6.81$ $50.67$ $32.5$ $8.07$ $36.43$ $65.53$ $65.05$ $60.16$ $73.29$ $63.85$
CD $6.81$ $51.07$ $31.87$ $8.66$ $35.58$ $66.92$ $64.65$ $60.22$ $73.23$ $63.14$
BCD(k=2) $6.65$ $50.86$ $31.67$ $8.91$ $35.15$ $67.21$ $64.4$ $60.31$ $73.45$ $63.22$
w4a16g128 OWC $7.12$ $47.85$ $33.58$ $9.99$ $35.58$ $67.09$ $68.38$ $60.15$ $73.5$ $62.9$
GPTQ $7.2$ $51.6$ $33.94$ $8.86$ $36.18$ $67.21$ $66.51$ $59.99$ $73.5$ $63.38$
CD $7.59$ $48.87$ $35.22$ $9.06$ $35.92$ $68.01$ $67.09$ $60.49$ $73.45$ $63.54$
BCD(k=2) $7.29$ $48.72$ $34.67$ $9.1$ $36.69$ $68.01$ $67.28$ $60.7$ $73.34$ $62.9$
OWC-CD $7.31$ $48.52$ $33.2$ $9.3$ $35.75$ $67.59$ $66.97$ $60.8$ $74.21$ $63.06$
OWC-CD + CD $7.15$ $49.15$ $34.77$ $9.4$ $36.52$ $67.8$ $66.33$ $60.74$ $73.61$ $63.61$
OWC-CD + BCD(k=2) $7.29$ $48.88$ $34.68$ $9.15$ $36.43$ $68.01$ $65.75$ $60.77$ $74.05$ $64.01$
Table 10: Table presents downstream evaluation (one-shot) numbers of PaLM2-Otter - for GPTQ, CD, BCD for INT2 quantization of FFN weights.
Method Epochs PaLM2-Otter
w2a16g128 NatualQ. SQuAD TriviaQA WebQ ARC-c ARC-e BoolQ HellaSwag PIQA WinoGrande
GPTQ - $6.93$ $64.89$ $28.36$ $7.73$ $44.28$ $71.04$ $78.81$ $68.29$ $72.69$ $67.72$
CD $1$ $8.23$ $68.12$ $32.98$ $10.14$ $43.69$ $72.43$ $76.91$ $68.11$ $72.96$ $66.46$
BCD(k=2) $1$ $8.09$ $69.2$ $32.85$ $9.99$ $43.43$ $73.02$ $77$ $68.03$ $73.07$ $66.38$
$2$ $8.2$ $69.43$ $32.98$ $9.74$ $43.52$ $73.7$ $77.03$ $68.56$ $74.1$ $67.8$
$3$ $8.31$ $69.79$ $33.03$ $9.65$ $43.52$ $73.11$ $77.09$ $68.76$ $73.07$ $67.01$
BCD(k=4) $1$ $7.87$ $69.52$ $33.24$ $9.79$ $43.17$ $72.6$ $77.83$ $68.61$ $73.88$ $66.46$
$2$ $8.31$ $69.06$ $33.03$ $9.6$ $43$ $72.6$ $76.64$ $68.51$ $73.18$ $66.3$
$3$ $8.31$ $69.8$ $32.69$ $9.5$ $42.15$ $72.43$ $77.06$ $68.34$ $73.72$ $67.88$
Table 11: Table presents downstream evaluation (one-shot) numbers of PaLM2-Bison - for GPTQ, CD, BCD for INT2 quantization of FFN weights.
Method Epochs PaLM2-Bison
w2a16g128 NatualQ. SQuAD TriviaQA WebQ ARC-c ARC-e BoolQ HellaSwag PIQA WinoGrande
GPTQ - $15.26$ $69.34$ $55.6$ $17.08$ $51.02$ $80.13$ $86.02$ $74.44$ $78.78$ $74.74$
CD $1$ $15.6$ $67.43$ $55.06$ $16.83$ $50.43$ $79.97$ $83.76$ $74.35$ $78.35$ $74.51$
BCD(k=2) $1$ $15.62$ $67.32$ $55.82$ $16.63$ $51.45$ $79.97$ $83.61$ $74.46$ $79$ $75.45$
$2$ $15.62$ $68.45$ $56.36$ $16.98$ $51.02$ $80.09$ $84.56$ $74.99$ $78.89$ $75.06$
$3$ $16.01$ $67.99$ $56.31$ $16.88$ $51.19$ $79.88$ $84.65$ $74.95$ $79.33$ $75.45$
BCD(k=4) $1$ $15.9$ $67.46$ $56.12$ $16.44$ $51.28$ $80.35$ $84.01$ $74.89$ $78.84$ $75.85$
$2$ $16.04$ $67.61$ $56.34$ $16.39$ $51.54$ $80.09$ $83.91$ $74.83$ $78.73$ $75.37$
$3$ $16.12$ $67.54$ $56.3$ $16.24$ $51.02$ $80.22$ $83.98$ $74.93$ $78.51$ $75.93$
Table 12: Table presents downstream evaluation (zero-shot) numbers for GPTQ, CD, BCD for INT3, INT4 quantization of FFN weights. Gen, Rank columns correspond to generation and ranking tasks.
Config Method PaLM2-Gecko PaLM2-Otter PaLM2-Bison
Gen. Rank Avg. Gen. Rank Avg. Gen. Rank Avg.
w16a16 - $25.85$ $61.44$ $47.2$ $44.31$ $75.43$ $62.98$ $51.08$ $78.28$ $68.00$
w3a16 OWC $15.45$ $55.73$ $39.62$ $15.06$ $59.29$ $41.60$ $39.58$ $75.45$ $61.10$
GPTQ $13.08$ $56.92$ $39.39$ $31.80$ $71.94$ $55.88$ $41.42$ $76.70$ $62.58$
CD $15.82$ $58.19$ $41.24$ $32.34$ $72.03$ $56.16$ $41.66$ $76.71$ $62.69$
BCD(k=2) $16.16$ $57.85$ $41.17$ $32.04$ $72.22$ $56.15$ $41.37$ $76.78$ $62.62$
w3a16g128 OWC $18.54$ $56.95$ $41.59$ $26.62$ $68.09$ $51.50$ $41.54$ $76.51$ $62.52$
GPTQ $17.96$ $58.30$ $42.16$ $33.43$ $72.98$ $57.16$ $43.47$ $76.65$ $63.38$
CD $16.81$ $58.10$ $41.58$ $32.73$ $72.62$ $56.66$ $42.95$ $77.03$ $63.4$
BCD(k=2) $16.33$ $57.81$ $41.22$ $32.13$ $72.67$ $56.46$ $42.67$ $76.85$ $63.18$
OWC-CD $16.73$ $57.74$ $41.34$ $32.90$ $72.40$ $56.60$ $42.23$ $77.02$ $63.11$
OWC-CD + CD $16.37$ $58.22$ $41.48$ $33.53$ $72.64$ $56.99$ $43.09$ $76.91$ $63.38$
OWC-CD + BCD(k=2) $16.47$ $58.45$ $41.66$ $33.71$ $72.73$ $57.12$ $43.07$ $76.70$ $63.25$
w4a16 OWC $16.07$ $59.42$ $42.08$ $33.76$ $72.22$ $56.83$ $43.59$ $77.13$ $63.71$
GPTQ $18.77$ $59.5$ $43.21$ $34.03$ $73.53$ $57.73$ $44.16$ $77.54$ $64.19$
CD $18.20$ $59.95$ $43.25$ $35.01$ $73.42$ $58.06$ $43.97$ $77.55$ $64.12$
BCD(k=2) $17.89$ $59.64$ $42.94$ $35.10$ $73.31$ $58.02$ $43.95$ $77.77$ $64.24$
w4a16g128 OWC $19.00$ $60.42$ $43.85$ $34.52$ $72.83$ $57.5$ $44.45$ $77.71$ $64.41$
GPTQ $19.07$ $60.40$ $43.87$ $35.34$ $73.60$ $58.29$ $44.25$ $77.65$ $64.29$
CD $20.35$ $60.94$ $44.70$ $36.20$ $73.67$ $58.68$ $44.21$ $77.72$ $64.31$
BCD(k=2) $20.44$ $60.69$ $44.59$ $36.11$ $73.54$ $58.57$ $44.17$ $77.56$ $64.2$
OWC-CD $19.80$ $60.39$ $44.15$ $36.33$ $73.62$ $58.71$ $44.75$ $77.70$ $64.52$
OWC-CD + CD $20.47$ $60.85$ $44.70$ $35.98$ $73.59$ $58.54$ $44.77$ $77.52$ $64.42$
OWC-CD + BCD(k=2) $20.15$ $60.49$ $44.36$ $36.00$ $73.62$ $58.57$ $44.70$ $77.70$ $64.50$
Table 13: Table presents downstream evaluation (zero-shot) numbers of PaLM2-Bison - for GPTQ, CD, BCD for INT3, INT4 quantization of FFN weights.
Config Method PaLM2-Bison
NatualQ. SQuAD TriviaQA WebQ ARC-c ARC-e BoolQ HellaSwag PIQA WinoGrande
w16a16 - $aa$ $aa$ $aa$ $aa$ $aa$ $aa$ $aa$ $aa$ $aa$ $aa$
w3a16 OWC $15.87$ $73.01$ $58.86$ $10.58$ $52.47$ $76.43$ $86.91$ $79.76$ $81.01$ $76.09$
GPTQ $17.31$ $71.25$ $65.53$ $11.56$ $54.44$ $79.25$ $87.22$ $80.78$ $81.61$ $76.87$
CD $17.84$ $71.24$ $64.76$ $12.8$ $54.44$ $79.17$ $87.58$ $80.79$ $82.1$ $76.16$
BCD(k=2) $17.81$ $71.3$ $64.29$ $12.06$ $53.5$ $79.08$ $87.65$ $80.91$ $82.21$ $77.35$
w3a16g128 OWC $15.98$ $71.28$ $68.02$ $10.88$ $53.58$ $79.55$ $86.73$ $80.9$ $81.61$ $76.72$
GPTQ $18.95$ $71.49$ $69.67$ $13.78$ $53.24$ $79.84$ $86.79$ $81.05$ $82.05$ $76.95$
CD $18.31$ $71.81$ $68.56$ $13.14$ $53.41$ $79.8$ $87.55$ $81.31$ $82.43$ $77.66$
BCD(k=2) $18.53$ $71.55$ $67.62$ $12.99$ $53.58$ $79.59$ $87.49$ $81.14$ $82.81$ $76.48$
OWC-CD $17.26$ $71.79$ $67.53$ $12.35$ $54.69$ $80.85$ $85.72$ $81.02$ $81.88$ $77.98$
OWC-CD + CD $18.75$ $72.05$ $68.35$ $13.19$ $54.1$ $80.18$ $86.36$ $80.86$ $82.48$ $77.51$
OWC-CD + BCD(k=2) $19.36$ $71.47$ $68.45$ $12.99$ $53.58$ $80.47$ $86.02$ $81.15$ $82.26$ $76.72$
w4a16 OWC $20$ $72.51$ $69.34$ $12.5$ $54.61$ $80.39$ $87.06$ $81.5$ $82.05$ $77.19$
GPTQ $21.25$ $72.39$ $70.6$ $12.4$ $55.2$ $80.64$ $87.58$ $81.98$ $82.26$ $77.58$
CD $21.5$ $71.57$ $70.04$ $12.8$ $55.29$ $80.56$ $88.07$ $81.78$ $82.59$ $77.03$
BCD(k=2) $21.16$ $71.92$ $70.01$ $12.7$ $55.63$ $81.19$ $87.77$ $81.82$ $82.81$ $77.43$
w4a16g128 OWC $21.22$ $73.06$ $70.07$ $13.44$ $55.8$ $80.47$ $87.77$ $82.08$ $82.64$ $77.51$
GPTQ $20.72$ $73.17$ $70.12$ $12.99$ $55.38$ $80.89$ $87.55$ $82.14$ $82.59$ $77.35$
CD $21.8$ $73.37$ $69.06$ $12.6$ $54.78$ $80.72$ $87.25$ $82.16$ $83.19$ $78.22$
BCD(k=2) $21.5$ $73.06$ $69.34$ $12.8$ $55.29$ $80.98$ $87.31$ $82.14$ $82.54$ $77.11$
OWC-CD $21.36$ $73.85$ $70.07$ $13.73$ $55.12$ $81.23$ $87.71$ $82.27$ $82.59$ $77.27$
OWC-CD + CD $21.41$ $73.32$ $71.46$ $12.89$ $54.18$ $81.4$ $87.28$ $82.17$ $82.37$ $77.74$
OWC-CD + BCD(k=2) $21.19$ $73.24$ $71.64$ $12.75$ $55.03$ $81.52$ $87.28$ $82.06$ $82.64$ $77.66$
Table 14: Table presents downstream evaluation (zero-shot) numbers of PaLM2-Otter - for GPTQ, CD, BCD for INT3, INT4 quantization of FFN weights.
Config Method PaLM2-Otter
NatualQ. SQuAD TriviaQA WebQ ARC-c ARC-e BoolQ HellaSwag PIQA WinoGrande
w16a16 - $aa$ $aa$ $aa$ $aa$ $aa$ $aa$ $aa$ $aa$ $aa$ $aa$
w3a16 OWC $2.27$ $49.45$ $8.02$ $0.49$ $41.98$ $60.14$ $48.78$ $66.91$ $71.87$ $66.06$
GPTQ $9.09$ $64.53$ $46.38$ $7.19$ $49.83$ $73.99$ $81.96$ $76.87$ $78.89$ $70.09$
CD $10.72$ $64.01$ $48.03$ $6.59$ $50.77$ $76.6$ $81.01$ $75.77$ $77.97$ $70.09$
BCD(k=2) $10.55$ $64.31$ $47.25$ $6.05$ $50.85$ $76.43$ $81.41$ $75.91$ $78.45$ $70.24$
w3a16g128 OWC $6.18$ $66.37$ $30.48$ $3.44$ $47.44$ $70.16$ $69.63$ $74.61$ $76.88$ $69.85$
GPTQ $10.17$ $66.53$ $49.29$ $7.73$ $51.37$ $75.93$ $82.35$ $77$ $79.11$ $72.14$
CD $7.17$ $68.73$ $49.42$ $5.61$ $50.26$ $76.18$ $82.39$ $76.79$ $78.13$ $71.98$
BCD(k=2) $6.68$ $68.18$ $48.39$ $5.27$ $50.26$ $75.93$ $81.41$ $76.84$ $79$ $72.61$
OWC-CD $8.98$ $68.37$ $48.33$ $5.91$ $50.51$ $75.46$ $81.35$ $76.83$ $78.18$ $72.06$
OWC-CD + CD $9.72$ $67.87$ $49.87$ $6.64$ $51.37$ $76.68$ $81.59$ $76.78$ $77.91$ $71.51$
OWC-CD + BCD(k=2) $9.94$ $68.26$ $49.91$ $6.74$ $50.85$ $76.22$ $81.68$ $76.99$ $78.89$ $71.74$
w4a16 OWC $11.19$ $66.67$ $49.56$ $7.63$ $50$ $73.91$ $80$ $77.36$ $79.11$ $72.93$
GPTQ $12.27$ $66.65$ $48.72$ $8.46$ $51.71$ $75.84$ $82.02$ $78.2$ $80.09$ $73.32$
CD $12.52$ $66.85$ $53.16$ $7.53$ $52.3$ $76.43$ $81.28$ $78.3$ $79.6$ $72.61$
BCD(k=2) $12.33$ $67$ $53.45$ $7.63$ $51.96$ $75.97$ $80.92$ $78.47$ $80.47$ $72.06$
w4a16g128 OWC $11.33$ $65.98$ $53.57$ $7.19$ $49.66$ $74.92$ $82.17$ $78.22$ $79.54$ $72.45$
GPTQ $11.5$ $65.81$ $56.36$ $7.68$ $51.71$ $76.52$ $82.08$ $78.86$ $79.27$ $73.16$
CD $11.72$ $66.85$ $57.63$ $8.61$ $51.19$ $76.6$ $82.51$ $78.89$ $79.49$ $73.32$
BCD(k=2) $11.66$ $66.73$ $57.5$ $8.56$ $51.19$ $76.3$ $82.29$ $78.88$ $79.27$ $73.32$
OWC-CD $12.27$ $66.28$ $58.21$ $8.56$ $51.62$ $76.85$ $82.32$ $78.85$ $79.49$ $72.61$
OWC-CD + CD $11.63$ $66.28$ $58.06$ $7.92$ $51.62$ $76.94$ $82.45$ $78.72$ $79.11$ $72.69$
OWC-CD + BCD(k=2) $11.72$ $65.96$ $58.29$ $8.02$ $51.79$ $76.77$ $82.17$ $78.8$ $79.49$ $72.69$
Table 15: Table presents downstream evaluation (zero-shot) numbers of PaLM2-Gecko - for GPTQ, CD, BCD for INT3, INT4 quantization of FFN weights.
Config Method PaLM2-Gecko
NatualQ. SQuAD TriviaQA WebQ ARC-c ARC-e BoolQ HellaSwag PIQA WinoGrande
w16a16 - $aa$ $aa$ $aa$ $aa$ $aa$ $aa$ $aa$ $aa$ $aa$ $aa$
w3a16 OWC $2.55$ $47.57$ $9.81$ $1.87$ $29.69$ $56.73$ $64.8$ $53.28$ $69.59$ $60.3$
GPTQ $3.35$ $44.13$ $3.97$ $0.89$ $30.2$ $62.12$ $63.79$ $56.33$ $71.22$ $57.85$
CD $3.55$ $45.71$ $11.32$ $2.71$ $30.63$ $60.56$ $66.45$ $57.2$ $71.87$ $62.43$
BCD(k=2) $3.63$ $47.04$ $11.22$ $2.76$ $30.29$ $60.52$ $65.32$ $57.33$ $71.82$ $61.8$
w3a16g128 OWC $3.68$ $55.38$ $12.5$ $2.61$ $31.66$ $60.14$ $66.12$ $53.19$ $71.71$ $58.88$
GPTQ $4.02$ $50.04$ $14.64$ $3.15$ $32.42$ $61.91$ $66.64$ $56.05$ $73.01$ $59.75$
CD $3.38$ $52.01$ $9.62$ $2.21$ $31.23$ $63.13$ $65.9$ $55.98$ $72.47$ $59.91$
BCD(k=2) $3.1$ $50.95$ $9.36$ $1.92$ $31.14$ $62.25$ $65.41$ $56.16$ $72.91$ $58.96$
OWC-CD $3.6$ $51.69$ $9.33$ $2.31$ $31.83$ $61.74$ $65.02$ $56.23$ $72.2$ $59.43$
OWC-CD + CD $3.82$ $50.39$ $9.5$ $1.77$ $32.51$ $62.25$ $66.33$ $56.91$ $71.27$ $60.06$
OWC-CD + BCD(k=2) $3.63$ $49.87$ $10.26$ $2.12$ $32.76$ $62.88$ $66.24$ $56.79$ $71.82$ $60.22$
w4a16 OWC $4.02$ $46.96$ $10.63$ $2.66$ $33.19$ $61.99$ $66.7$ $59.26$ $73.67$ $61.72$
GPTQ $4.46$ $48.69$ $18.84$ $3.1$ $34.47$ $62.96$ $64.77$ $60.09$ $73.61$ $61.09$
CD $4.13$ $48.88$ $16.64$ $3.15$ $33.53$ $63.89$ $66.79$ $60.31$ $73.78$ $61.4$
BCD(k=2) $4.24$ $48.31$ $15.73$ $3.3$ $33.28$ $63.34$ $66.39$ $60.26$ $73.56$ $61.01$
w4a16g128 OWC $4.99$ $46.1$ $20.62$ $4.28$ $34.73$ $64.48$ $67.83$ $59.91$ $74.32$ $61.25$
GPTQ $5.18$ $49.79$ $18.17$ $3.15$ $35.15$ $64.6$ $66.24$ $60.21$ $74.32$ $61.88$
CD $5.73$ $46.96$ $24.17$ $4.53$ $36.18$ $66.04$ $67.8$ $60.7$ $73.67$ $61.25$
BCD(k=2) $5.79$ $47.28$ $24.11$ $4.58$ $36.01$ $65.7$ $67.06$ $60.73$ $73.61$ $61.01$
OWC-CD $5.54$ $47.24$ $21.71$ $4.72$ $34.3$ $64.73$ $68.01$ $60.72$ $73.56$ $61.01$
OWC-CD + CD $5.62$ $47.6$ $24$ $4.68$ $35.49$ $66.16$ $67.06$ $60.78$ $74.1$ $61.48$
OWC-CD + BCD(k=2) $5.57$ $47.17$ $23.05$ $4.82$ $35.24$ $65.82$ $66.02$ $60.9$ $73.23$ $61.72$
Table 16: Table presents downstream evaluation (zero-shot) numbers for GPTQ, CD, BCD for INT2 quantization of FFN weights. Gen, Rank columns correspond to generation and ranking tasks.
Method Epochs PaLM2-Otter PaLM2-Bison
w2a16g128 Generation Rank Avg. Generation Rank Avg.
GPTQ - $23.06$ $66.62$ $49.2$ $33.62$ $72.74$ $57.09$
CD 1 $25.12$ $65.73$ $49.48$ $33.63$ $71.80$ $56.53$
BCD(k=2) $1$ $25.42$ $65.57$ $49.51$ $33.97$ $72.16$ $56.88$
$2$ $24.73$ $65.78$ $49.36$ $34.01$ $72.39$ $57.04$
$3$ $25.28$ $66.01$ $49.72$ $34.27$ $72.67$ $57.31$
BCD(k=4) $1$ $25.15$ $66.04$ $49.68$ $33.94$ $72.10$ $56.84$
$2$ $24.86$ $65.71$ $49.37$ $33.87$ $71.97$ $56.73$
$3$ $25.70$ $65.30$ $49.46$ $33.85$ $72.06$ $56.78$
Table 17: Table presents downstream evaluation (zero-shot) numbers of PaLM2-Otter - for GPTQ, CD, BCD for INT2 quantization of FFN weights.
Method Epochs PaLM2-Otter
w2a16g128 NatualQ. SQuAD TriviaQA WebQ ARC-c ARC-e BoolQ HellaSwag PIQA WinoGrande
GPTQ - $4.32$ $63.43$ $22.08$ $2.41$ $42.32$ $66.96$ $80.09$ $68.54$ $73.78$ $68.03$
CD $1$ $4.99$ $66.51$ $25.53$ $3.44$ $40.87$ $70.33$ $74.89$ $68.13$ $72.96$ $67.17$
BCD(k=2) $1$ $4.93$ $67.51$ $25.56$ $3.69$ $41.47$ $69.7$ $74.98$ $67.97$ $73.72$ $65.59$
$2$ $4.6$ $67.74$ $23.6$ $3$ $40.78$ $69.49$ $75.47$ $68.27$ $73.39$ $67.25$
$3$ $5.1$ $67.91$ $24.53$ $3.59$ $42.41$ $69.87$ $75.02$ $68.52$ $73.88$ $66.38$
BCD(k=4) $1$ $4.9$ $67.5$ $25.01$ $3.2$ $42.15$ $68.69$ $76.18$ $68.51$ $73.61$ $67.09$
$2$ $4.79$ $67.9$ $23.55$ $3.2$ $41.81$ $69.82$ $73.39$ $68.5$ $73.5$ $67.25$
$3$ $5.46$ $68.29$ $26$ $3.05$ $41.21$ $69.02$ $73.98$ $68.06$ $72.52$ $67.01$
Table 18: Table presents downstream evaluation (zero-shot) numbers of PaLM2-Bison - for GPTQ, CD, BCD for INT2 quantization of FFN weights.
Method Epochs PaLM2-Bison
w2a16g128 NatualQ. SQuAD TriviaQA WebQ ARC-c ARC-e BoolQ HellaSwag PIQA WinoGrande
GPTQ - $11.55$ $66.72$ $48.79$ $7.43$ $47.27$ $76.89$ $84.56$ $74.36$ $79.54$ $73.8$
CD $1$ $11.25$ $64.49$ $50.84$ $7.92$ $48.12$ $75.97$ $79.39$ $74.28$ $79.11$ $73.95$
BCD(k=2) $1$ $11.77$ $64.17$ $51.36$ $8.56$ $48.38$ $75.88$ $81.01$ $74.6$ $78.84$ $74.27$
$2$ $11.61$ $64.69$ $51.23$ $8.51$ $47.87$ $76.09$ $81.74$ $74.72$ $79$ $74.9$
$3$ $11.72$ $64.5$ $52.15$ $8.71$ $48.46$ $76.18$ $82.54$ $74.82$ $79.6$ $74.43$
BCD(k=4) $1$ $11.69$ $64.48$ $51.07$ $8.51$ $47.7$ $75.63$ $80.89$ $74.55$ $79.6$ $74.27$
$2$ $11.75$ $64.39$ $51.36$ $7.97$ $48.46$ $75.04$ $80.58$ $74.48$ $79.38$ $73.88$
$3$ $11.8$ $64.41$ $51.16$ $8.02$ $48.12$ $75.55$ $80.86$ $74.55$ $78.73$ $74.59$
Appendix E Quantization Runtime
Table 19: Table presents the time taken (in minutes) to quantize PaLM2-Otter and PaLM2-Bison. All models were quantized with 8 80GB H100s, and run for 1 epoch (i.e., $T=d_{\text{in}}$).
Config Method PaLM2-Otter PaLM2-Bison
FFN runtime
(in minutes)
Attn. runtime
(in minutes)
FFN runtime
(in minutes)
Attn. runtime
(in minutes)
w3a16 GPTQ $0.90$m $0.72$m $6.42$m $0.63$m
CD $2.94$m $0.49$m $16.13$m $0.43$m
BCD(k=2) $14.03$m $2.08$m $82.13$m $1.81$m
w3a16g128 GPTQ $1.02$m $0.75$m $6.70$m $0.66$m
CD $3.32$m $0.59$m $18.44$m $0.52$m
BCD(k=2) $14.73$m $2.12$m $88.10$m $1.86$m
w4a16 GPTQ $0.90$m $0.64$m $6.60$m $0.56$m
CD $3.59$m $0.59$m $21.19$m $0.52$m
BCD(k=2) $34.85$m $4.72$m $235.14$m $4.13$m
w4a16g128 GPTQ $0.98$m $0.69$m $6.962$m $0.60$m
CD $3.99$m $0.69$m $23.48$m $0.60$m
BCD(k=2) $42.37$m $5.41$m $192.59$m $4.74$m
Appendix F Comparison to QuantEase
We also compare our per-channel quantization results with those of QuantEase, a parallel study to ours. It’s important to note that QuantEase does not have a publicly available implementation. So, we implemented the cyclic coordinate descent strategy used by QuantEase, and used the same initialization and regularization strength as our algorithms (although the QuantEase paper doesn’t provide these details). We then ran QuantEase for the recommended number of iterations specified in the paper ($20$ epochs or $T=20d_{\text{in}}$ iterations). The findings are presented in Table 20. On PaLM2-Gecko, PaLM2-Otter, and PaLM2-Bison, both these approaches have similar performance. These results collectively highlight the potential of coordinate descent algorithms in outperforming GPTQ.
Config Method PaLM2-Gecko PaLM2-Otter PaLM2-Bison
w3a16 GPTQ $11.347$ $7.176$ $5.774$
QuantEase (epochs=20) $10.731$ $6.996$ $5.741$
CD $10.920$ $7.002$ $5.739$
BCD(k=2) $10.898$ $6.979$ $5.733$
w4a16 GPTQ $8.764$ $6.249$ $5.417$
QuantEase (epochs=20) $8.670$ $6.197$ $5.408$
CD $8.694$ $6.195$ $5.407$
BCD(k=2) $8.691$ $6.192$ $5.405$
Table 20: Comparison of QuantEase with CD and BCD for FFN quantization.